OpenStudy (anonymous):

What is the solution to this equation? 5x - 4 + 3x = 36 A. x = 4 B. x = 5 C. x = 20 D. x = 16

OpenStudy (anonymous):

Hey daijagattison!!! What about to start thinking??? We are not machine for automatic HW doing....

OpenStudy (anonymous):

A is the answer

OpenStudy (anonymous):

-3x+4y=1 -6x+6y=-3 Solve using the elimination method. show your work. If the system has no solution or an infinite number of solutions, state this. can someone please help

OpenStudy (anonymous):

I need help with this one first

OpenStudy (anonymous):

where is the ask a question field located

OpenStudy (anonymous):

DID YOU GET IT

OpenStudy (anonymous):

SORRY I DO NOT SEE WHERE IT IS

OpenStudy (anonymous):

i HAVE BEEN DOING IT THIS WAY WITH NO PROBLEM

OpenStudy (anonymous):

Please stop spamming in this question...

OpenStudy (anonymous):

WHERE DO I POST MY PROBLEM THEN?

OpenStudy (anonymous):

Hi @connielynn - I see you're having some difficulty, hopefully I can help. Thank you @matt1888 for your help btw, cheers ;). So, @connielynn when you want to ask a question, you should see a box that says 'Ask a question...' in the upper left. Should be right under the blue bar that says 'Mathematics' - are you able to find it?

OpenStudy (anonymous):

DO NOT SEE ANYTHING I SEE THE BLUE BAR AT THE TOP LEFT HAND CORNER THAT SAYS MATHMATICS, AND UNDER THAT IS A GRAY BAR AND IT SAYS YOU HAVE AN OPEN QUESTION. SORRY

OpenStudy (anonymous):

Hi @connielynn - no problem, also, please take off the CAPS LOCK (it's kinda like yelling online). So, you have an open question, if you've gotten the answer you need for that question, then simply 'Close' the question and you can ask another.